OpenStudy (anonymous):

find the solution for the logarithmic equation ln(x+8) +ln(x-8)=0

OpenStudy (anonymous):

x=-7 and x=9

OpenStudy (anonymous):

(x+8)(x-8)=1 Solve th quadratic

OpenStudy (anonymous):

-7 is not a valid solution

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

Log(x-8)=0

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

you cannot define logarithm for a negative number

OpenStudy (anonymous):

You can

OpenStudy (anonymous):

i dont think she knows complex logarithms

OpenStudy (anonymous):

yes you can say log (17e^(i*pi)) .....but thats not what she might want

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=ln%28x%2B8%29+%2Bln%28x-8%29%3D0 This is how I would look at this one. :)

OpenStudy (anonymous):

It can be negative

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

Look it up if ya like.

OpenStudy (anonymous):

wolfram shows were both fools....root65 is the answer

OpenStudy (anonymous):

i told you i know there can be negative logarithms...they just dont need em at this level...there are even logs of quantities that are sums of scalars and vectors..would you want to know that before your exam?

OpenStudy (anonymous):

i don't get it, can someone show their work?

OpenStudy (anonymous):

ln(x+8) +ln(x-8)=0 ln a + lnb = ln(a.b) (a +b) ( a - b) = a^2 - b^2 (x+8) (x-8) = ( x^2 - 8^2) -> ln (x^2 - 64) = 0 Exponent both side by e: => x^2 - 64 = e^0 = 1 => x^2 = 64 + 1 = 65 Thus x = +/- sqrt (65) = 8.06, -8.06